Problem M6-01


You are an industry analyst who specializes in an industry where the market inverse demand is P = 200 - 4Q. The external marginal cost of producing the product is MCExternal= 6Q, and the internal cost is MCInternal = 12Q.

a. What is the socially efficient level of output?

b. Given these costs and market demand, how much output would a competitive industry produce?

c. Given these costs and market demand, how much output would a monopolist produce?

d. Discuss actions the government might take to induce firms in this industry
to produce the socially efficient level of output?
Answer #1

a. Socially efficient level of output?

set P = MCexternal + MCinternal to get

200 - 4Q = 18Q
200 = 22Q
Q = 9.1 units.

b. Output would a competitive industry produce:

Set P = MCinternal to get

200 - 4Q = 12Q
200 = 16Q
Q = 12.5 units

c. The output for monopolist produce:

set MR = MCinternal
200 - 8Q = 12Q
200 = 20Q
Q = 10 units.

d. Pollution taxes or permits can help reduce firms to produce the socially efficient level of output.

answered by: Aarti Saxena